If a car at rest accelerates uniformly to a speed of 144 km/hr in 20seconds then it covers distance of

Grade:11

1 Answers

Utsav Basu
6 years ago
u = 0
v = 144km/hr = 40m/s
t = 20s
Using, v = u + at, we get
40 = a*20
or, a = 2
Now, using s = (at2)/2, we get
s = (2*(20)2)/2
   = 400
Thus, distance covered is 400m